OpenStudy (anonymous):

Factor completely: 2x^3 + 4x^2 + 6x + 12 A.2[(x^2 + 3)(x + 2)] B.(2x^2 + 6)(x + 2) C.2[(x^2 + 2)(x + 3)] D.Prime

OpenStudy (mertsj):

Factor by grouping. Factor 2x^2 out of the first two terms. Factor 6 out of the last two terms. Then you will see that x+2 is a common factor. Factor.

Directrix (directrix):

Following @Mertsj instructions: 2x^3 + 4x^2 + 6x + 12 = 2x^2( x + 2) + 6(x + 2) Note that (x+2) is a common factor.

Directrix (directrix):

So, 2x^2( x + 2) + 6(x + 2) = ( x + 2)* ( ? + ? ) @isabellaxo